Q=6L^2-0.4L^3 is the function of short run production
Find the value of L that maximumes the out put
1
Expert's answer
2020-06-10T19:12:45-0400

We have the production function as:



"Q=6L^2-0.4L^3"

From the production function, the marginal product of labor is:



"MPL = 12L - 0.12L^2"

At the optimal production, the marginal product is zero. Therefore:



"MPL = 12L - 0.12L^2 = 0\\[0.3cm]
12L - 0.12L^2 = 0\\[0.3cm]
12 = 0.12L\\[0.3cm]
L = \dfrac{12}{0.12}\\[0.3cm]
\color{red}{L^* = 100}"